Hi,
This math problem confuses me very much....
Bill is three times older than Susan. In 18 years Bill is only 2 times older than Susan. Find the present age of Bill and Susan.
Thanks

Answer by ankor@dixie-net.com(22740)

An easy to understand method
Let b = Bill' present age
Let s = Susan's age
:
Write an equation for each statement
:
"Bill is three times older than Susan."
b = 3s
"In 18 years Bill is only 2 times older than Susan."
Both will age 18 yrs, so write the equation:
b + 18 = 2(s+18)
b + 18 = 2s + 36
b = 2s + 36 - 18
b = 2s + 18
From the 1st statement we know, that b = 3s, therefore
3s = 2s + 18
3s - 2s = 18
s = 18 yrs is Susan's age
then obviously;
3(18) = 54 yrs Bill's age
;
:
Check this in the 2nd statement
54 + 18 = 2(18 + 18)
72 = 2(36)